Use the information below to answer the next question

52W    52W                                           Yield                       Vol                                                         Net

Hi         lo Ticker     Div                %            P/E          00s        High        Low        Close       Chg

43.21   24.68     PG        1.17               3.0          25.46    6412      40.02     38.37        ?             -0.63

What is the lowest the stock has sold for today?

What is the closing price of the stock?

What was the previous day’s closing price?

What is the EPS for the stock?
Answer #1
Lowest stock price today $ 38.37
Closing price $ 39.00 =1.17/3%
Previous day’s closing price $ 39.63 =39+0.63
EPS $    1.53 =39/25.46